OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3).

OpenStudy (anonymous):

f(3)=3^2-1=8

OpenStudy (anonymous):

substitute .............

OpenStudy (anonymous):

x squared-1?

OpenStudy (anonymous):



OpenStudy (anonymous):

so its 8

OpenStudy (anonymous):

=(3)^2 -1 =9-1 =8 :)